# Caption-Driven Explainability: Probing CNNs for Bias via CLIP

### v1

@@ Abstract
Robustness has become one of the most critical problems in machine learning (ML). The science of interpreting ML models to understand their behavior and improve their robustness is referred to as explainable artificial intelligence (XAI). One of the state-of-the-art XAI methods for computer vision problems is to generate saliency maps. A saliency map highlights the pixel space of an image that excites the ML model the most. However, this property could be misleading if spurious and salient features are present in overlapping pixel spaces. In this paper, we propose a caption-based XAI method, which integrates a standalone model to be explained into the contrastive language-image pre-training (CLIP) model using a novel network surgery approach. The resulting caption-based XAI model identifies the dominant concept that contributes the most to the models prediction. This explanation minimizes the risk of the standalone model falling for a covariate shift and contributes significantly towards developing robust ML models.
@@ 1 Introduction
The fundamental property of ML models is that they are not explicitly programmed but learn from data instead. This attribute makes advanced ML models very powerful but challenging to interpret. With the ever-increasing capabilities and importance of ML models at the core of many applications, there is a need to prove their robustness, which represents one of the most crucial research areas in artificial intelligence (AI).[1] A robust ML model’s performance in real-world situations deviates only marginally from the test performance, even if one or more features change drastically due to unforeseen circumstances. Expressing it differently, robustness refers to a model’s ability to resist being fooled. In theory, the training, validation, and test datasets are sampled from the same data distribution. The temptation to deploy a low bias, low variance ML model as shown in Fig. 1 is high. In a real-world scenario, there is always a risk involved that the data used for the training, validation, and test datasets does not accurately reflect the data distribution the deployed model faces. This distribution shift between the data used during the development of the model and the deployed model is designated as a covariate shift [2]. A covariate shift may be responsible for a model working in the lab for its intended task while failing in the real world. This characteristic is especially challenging in high-stakes environments, e.g., in medicine, where a patient could suffer from incorrect predictions made by an ML model [3]. One evident approach to avoid a covariate shift is to ensure that the data for the development of the model reflects the real-world perfectly, but this is by no means a trivial task. Another approach is to use XAI methods. One of the state-of-the-art XAI methods to improve the robustness of ML models for computer vision problems is to generate saliency maps. There is a large variety of possibilities to generate saliency maps using class activation maps (CAM) [4], gradient-weighted CAM (Grad-CAM) [5], or learning important features CAM (LIFT-CAM) [6], which estimates shapley values to weight the linear combination of activation maps by their marginal contribution to the explanation. All saliency map methods highlight the pixel space of an image that excites the model the most. [7] However, this property could be misleading if spurious and salient features are present in overlapping pixel spaces [8]. The work of Bau et al. [9] about the GAN dissection method suggests that it is essential to understand the internal concepts of a model since these insights can help to improve the model’s behavior. Their network dissection method [10] demonstrates the generalizability of individual units responding to specific high-level concepts not directly represented in the training dataset. Measuring the alignment between the unit response and a set of concepts drawn from the broad and dense segmentation dataset [11] enables to define units as specific concept detectors. Inspired by the work on discovering concepts by Bau et al., our method, the caption-based XAI method, incorporates text to enhance the explanation. The main contribution of this paper solves the problem of identifying the dominant concept in multimodal units and therefore revealing a potential covariate shift before the deployment of the standalone model. Additionally, the caption-based XAI method works reliably even if spurious and salient features are present in overlapping pixel spaces. The demonstration of the caption-based XAI method in this paper uses a biased dataset, which leads to a biased standalone model. The biased dataset contains a covariate shift between the train, validation, and test datasets (representing the available data during the model development) and the real-world dataset (representing real-world data after deployment). The objective of the standalone model is to classify handwritten digits with the values five and eight from the MNIST dataset [12]. In the data available during the model development, all digits with the value five are colored red and all digits with the value eight are colored green. In the real-world dataset, the color assignments are random. This difference in the color assignment is responsible for the covariate shift. The caption-based XAI method aims to identify the dominant concept of the standalone model.
@@ 2 Proposed method
The proposed caption-based XAI method uses a network surgery process to transfer the properties from the standalone model to be explained into CLIP [13] by swapping similar activation maps from the standalone image encoder to the CLIP image encoder resulting in the caption-based XAI model as shown in Fig. 2. Derived from the Euclidean dot product, the cosine similarity denotes the alignment of the two embeddings $\boldsymbol{I}_{i}$ and $\boldsymbol{T}_{j}$ in CLIP’s space of concepts.
@@
Using suitable captions, the texts describing dominant concepts in the images result in significant embedding similarities. [14] If these high scores primarily arise for the color descriptions, then the standalone model is color biased. If these high scores primarily arise for the shape descriptions, then the standalone model is focused on the shapes.
@@ 2.1 Architecture
CLIP is the core component of the caption-based XAI model. Many different configurations are available for CLIP’s text and image encoder. Throughout this paper, the CLIP text encoder is a masked self-attention transformer [15, 16] and the CLIP image encoder is OpenAI’s modified and pre-trained [17] residual neural network-52 (ResNet-52) [18] model. The main modifications of the ResNet model are the addition of two convolutional layers in the first stage and the replacement of the average pooling layer with an attention pooling layer. There are 51 convolutional layers, one fully connected layer and two pooling layers in the CLIP image encoder.
The standalone image encoder to be explained is a ResNet-50 model, which has been pre-trained on the ImageNet dataset [19] and finetuned for the MNIST binary classification task. There are 49 convolutional layers, one fully connected layer and two pooling layers in the standalone model.

### v4

@@ Abstract
Robustness has become one of the most critical problems in machine learning (ML). The science of interpreting ML models to understand their behavior and improve their robustness is referred to as explainable artificial intelligence (XAI). One of the state-of-the-art XAI methods for computer vision problems is to generate saliency maps. A saliency map highlights the pixel space of an image that excites the ML model the most. However, this property could be misleading if spurious and salient features are present in overlapping pixel spaces. In this paper, we propose a caption-based XAI method, which integrates a standalone model to be explained into the contrastive language-image pre-training (CLIP) model using a novel network surgery approach. The resulting caption-based XAI model identifies the dominant concept that contributes the most to the models prediction. This explanation minimizes the risk of the standalone model falling for a covariate shift and contributes significantly towards developing robust ML models. Our code is available at https://github.com/patch0816/caption-driven-xai.
@@ 1 Introduction
The fundamental property of ML models is that they are not explicitly programmed but learn from data instead. This attribute makes advanced ML models very powerful but challenging to interpret. With the ever-increasing capabilities and importance of ML models at the core of many applications, there is a need to prove their robustness, which represents one of the most crucial research areas in artificial intelligence (AI).[adadi2018peeking] A robust ML model’s performance in real-world situations deviates only marginally from the test performance, even if one or more features change drastically due to unforeseen circumstances. Expressing it differently, robustness refers to a model’s ability to resist being fooled. In theory, the training, validation, and test datasets are sampled from the same data distribution. The temptation to deploy a low bias, low variance ML model as shown in Fig. 1 is high. In a real-world scenario, there is always a risk involved that the data used for the training, validation, and test datasets does not accurately reflect the data distribution the deployed model faces. This distribution shift between the data used during the development of the model and the deployed model is designated as a covariate shift [neuhaus2023spurious]. A covariate shift may be responsible for a model working in the lab for its intended task while failing in the real world. This characteristic is especially challenging in high-stakes environments, e.g., in medicine, where a patient could suffer from incorrect predictions made by an ML model [degrave2021ai]. One evident approach to avoid a covariate shift is to ensure that the data for the development of the model reflects the real-world perfectly, but this is by no means a trivial task. Another approach is to use XAI methods. One of the state-of-the-art XAI methods to improve the robustness of ML models for computer vision problems is to generate saliency maps. There is a large variety of possibilities to generate saliency maps using class activation maps (CAM) [zhou2016learning], gradient-weighted CAM (Grad-CAM) [selvaraju2017grad], or learning important features CAM (LIFT-CAM) [jung2021towards], which estimates shapley values to weight the linear combination of activation maps by their marginal contribution to the explanation. All saliency map methods highlight the pixel space of an image that excites the model the most. [Ahn_2024_CVPR] However, this property could be misleading if spurious and salient features are present in overlapping pixel spaces [rudin2019stop]. The work of Bau et al. [bau2019gandissect] about the GAN dissection method suggests that it is essential to understand the internal concepts of a model since these insights can help to improve the model’s behavior. Their network dissection method [bau2020units] demonstrates the generalizability of individual units responding to specific high-level concepts not directly represented in the training dataset. Measuring the alignment between the unit response and a set of concepts drawn from the broad and dense segmentation dataset [oikarinen2023clipdissectautomaticdescriptionneuron] enables to define units as specific concept detectors. Inspired by the work on discovering concepts by Bau et al., our method, the caption-based XAI method, incorporates text to enhance the explanation. The main contribution of this paper solves the problem of identifying the dominant concept in multimodal units and therefore revealing a potential covariate shift before the deployment of the standalone model. Additionally, the caption-based XAI method works reliably even if spurious and salient features are present in overlapping pixel spaces. The demonstration of the caption-based XAI method in this paper uses a biased dataset, which leads to a biased standalone model. The biased dataset contains a covariate shift between the train, validation, and test datasets (representing the available data during the model development) and the real-world dataset (representing real-world data after deployment). The objective of the standalone model is to classify handwritten digits with the values five and eight from the MNIST dataset [lecun1998mnist]. In the data available during the model development, all digits with the value five are colored red and all digits with the value eight are colored green. In the real-world dataset, the color assignments are random. This difference in the color assignment is responsible for the covariate shift. The caption-based XAI method aims to identify the dominant concept of the standalone model.
@@ 2 Proposed method
The proposed caption-based XAI method uses a network surgery process to transfer the properties from the standalone model to be explained into CLIP [radford2021learning] by swapping similar activation maps from the standalone image encoder to the CLIP image encoder resulting in the caption-based XAI model as shown in Fig. 2. Derived from the Euclidean dot product, the cosine similarity denotes the alignment of the two embeddings $\boldsymbol{I}_{i}$ and $\boldsymbol{T}_{j}$ in CLIP’s space of concepts.
@@
Using suitable captions, the texts describing dominant concepts in the images result in significant embedding similarities. [bhalla2024splice] If these high scores primarily arise for the color descriptions, then the standalone model is color biased. If these high scores primarily arise for the shape descriptions, then the standalone model is focused on the shapes.
@@ 2.1 Architecture
CLIP is the core component of the caption-based XAI model. Many different configurations are available for CLIP’s text and image encoder. Throughout this paper, the CLIP text encoder is a masked self-attention transformer [vaswani2017attention, radford2019language] and the CLIP image encoder is OpenAI’s modified and pre-trained [zhang2019making] residual neural network-52 (ResNet-52) [he2016deep] model. The main modifications of the ResNet model are the addition of two convolutional layers in the first stage and the replacement of the average pooling layer with an attention pooling layer. There are 51 convolutional layers, one fully connected layer and two pooling layers in the CLIP image encoder.
The standalone image encoder to be explained is a ResNet-50 model, which has been pre-trained on the ImageNet dataset [imagenet_challenge] and finetuned for the MNIST binary classification task. There are 49 convolutional layers, one fully connected layer and two pooling layers in the standalone model.